OpenStudy (anonymous):

a disc jockey must play 12 commercial spots during 1 hour of radio show. each commercial is either 30 or 60 seconds long. if the total commercial time during 1 hour is 8 minutes, how many 30 second commercials were played that hour? how many 60 second commercials?

OpenStudy (anonymous):

assume there were a 30-second commercials assume there were b 60-second commercials then: a + b = 12 30*a + 60*b = 8*60 tidy up the second equation: a + 2b = 16 use elimination (equation 2 - equation 1): b = 4 substitute back in: a = 8 so there were 8 30-second commercials and there were 4 60-second commercials